# Tera-MIND: Tera-scale mouse brain simulation via spatial mRNA-guided diffusion

### v2

@@ 2.3.2 DOPA neuronal system
In dopamine-mediated neuronal signaling, Nr4a2 (Nurr1) is an important transcription factor that is expressed in several regions of the central nervous system including the substantia nigra (SN), ventral tegmental area (VTA), and olfactory bulb [16]. Nr4a2 plays a pivotal role in dopaminergic neurons by regulating several key genes (e.g., Th) involved in dopamine synthesis, storage, and release. Th encodes tyrosine hydroxylase, the rate-limiting enzyme in dopamine synthesis [16]. Nr4a2 and Th expression have been shown to co-localize in areas such as SN, VTA, and olfactory bulb of the mouse brain [33, 13]. Such co-localization patterns have been correctly captured in spatial visualizations presented in Fig. 3 (e, f). Reduced Nr4a2 and Th expression have been associated with Parkinson’s disease [18, 11], a common neurodegenerative disorder with loss of dopaminergic neurons in the SN, leading to dopamine deficiency and motor function impairments [18, 11]. This has also been observed in mice after deletion of Nr4a2 in mature dopaminergic neurons [10]. In the analysis of DOPA neuronal signaling, Tera-MIND uncovers distinct and strong Nr4a2-Th attention signals in the SN, VTA, and olfactory bulb, where both genes play a central role in regulating dopaminergic functions. As shown in Fig. 3 (e, g, h), the 3D gene-gene attention map reveals a spatially aligned interaction between Nr4a2 and Th that reinforces their roles in the molecular regulation of dopaminergic pathways and cellular anatomical structures.
@@
Figure 3: The visualization of gene-gene interactions for GLUT and DOPA neuronal systems (Main result). a. The 3D visualization of Slc17a6 and Slc17a7 gene expression (left), the conceptual illustration of 3D gene-gene attention layer (middle), and the 3D visualization of Slc17a6-Slc17a7 attention (right). b. The WSI visualization of Slc17a6 and Slc17a7 gene expression (expr.). c. The trending line showing the relationship between gene expr. and attention level for both Slc17a6 and Slc17a7. d. The WSI visualization of Slc17a6-Slc17a7 attention level. e. The 3D visualization of Nr4a2 and Th gene expression (left), the conceptual illustration of 3D gene-gene attention layer (middle), and the 3D visualization of Nr4a2-Th attention (right). f. The WSI visualization of Nr4a2 and Th gene expression (expr.). g. The trending line showing the relationship between gene expr. versus attention level for both Nr4a2 and Th. h. The WSI visualization of Nr4a2-Th attention level. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for both GLUT and DOPA pathways.
@@
2.4 Tera-MIND achieves reproducible and robust simulation results on three tera-scale virtual mouse brains.
In addition to the main mouse brain discussed above, two additional P56 adult mouse brains – one male and one female – were profiled in previous studies [47, 49] for supporting analyses, as illustrated in Fig. 4 (a) (middle and right). To rigorously evaluate the reproducibility and robustness of our findings, we extend the same simulation experiments conducted on the main mouse brain to both supporting mouse brains (supp (m) and supp (f)).
Morphology generation: As illustrated in Fig. 4 (a, b), the side-by-side visual comparisons between our generated and GT results demonstrate convincing reconstructions of brain morphology on all three tera-scale mouse brains. Consistent with the main generation results (Fig. 1 (b) and the left plots of Fig. 4 (a, b)), our supporting experiments highlight competitive and reliable 3D and slice-wise image generation quality, as elaborated by the holistic and multi-scale visualization shown in the middle and right plots of Fig. 4 (a, b). For quantitative evaluation, Tera-MIND is benchmarked against seven SOTA methods COCO-GAN [22], InfinityGAN [23], MS-PIE [46], SST-editing [44], IST-editing [42], SinFusion [28], and Patch-DM [12]. Complementary to PSNR and SSIM scores, we incorporate the spatial Fréchet Inception Distance (sFID) [26], which prioritizes spatially relevant evaluations by rewarding image distributions that preserve coherent spatial structures. Across all three brain instances, Tera-MIND consistently outperforms competing methods (Fig. 4), in terms of achieving better PSNR, SSIM, and sFID scores. When evaluated on domain-specific metrics such as cell number and nuclear size, Tera-MIND yields minimal distributional discrepancies relative to GT. This highlights its capacity to maintain high-fidelity not only in structural reconstruction but in the reproduction of biologically relevant features.
Pathway identification: In the context of the GLUT pathway, Fig. 5 (a, b) present a highly consistent and spatially comparable localization of Slc17a6 and Slc17a7 expression patterns across two independent supporting instances. These expression maps are accurately registered in 3D space and align well with the ones from our main results and previous studies [47, 49]. Similarly as observed in the main results, Fig. 5 (c) and (d) illustrate that Tera-MIND is capable of learning strong and heterogeneous Slc17a6-Slc17a7 attention levels driven by the positively correlated gene expression levels. This suggests their synergistic and collective roles in regulating neurotransmission and cellular morphology across overlapped regions. Regarding the DOPA pathway, despite the absence of the olfactory bulb in the raw data of the supporting mouse brains, the gene expression patterns of Nr4a2 and Th exhibit a comparable emergence in the SN and VTA regions (See also Fig. 5 (e, f)). Consistent with the main results, we observe reproducible and clear identification of Nr4a2-Th attention levels derived from Tera-MIND in the same regions for both supporting mouse brains (Fig.5 (e, h)). These results confirm a consistent expression pattern of critical genes associated with dopaminergic function in these specific regions. For more comprehensive 3D illustrations of both pathways, we refer interested readers to the side-by-side video visualizations across all three mouse brain instances.
@@
Figure 5: The visualization of gene-gene interaction for GLUT and DOPA neuronal systems (Supp (m) and (f) results). a. The 3D visualization of Slc17a6 and Slc17a7 gene expression (left) and the 3D visualization of Slc17a6-Slc17a7 attention (right). b. The WSI visualization of Slc17a6 and Slc17a7 gene expression (expr.). c. The trending line showing the relationship between gene expr. and attention level for both Slc17a6 and Slc17a7. d. The WSI visualization of Slc17a6-Slc17a7 attention level. e. The 3D visualization of Nr4a2 and Th gene expression (left) and the 3D visualization of Nr4a2-Th attention (right). f. The WSI visualization of Nr4a2 and Th gene expression (expr.). g. The trending line showing the relationship between gene expr. and attention level for both Nr4a2 and Th. h. The WSI visualization of Nr4a2-Th attention level. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for both GLUT and DOPA pathways.
3 Discussion
In this study, we introduced a novel patch-based diffusion model Tera-MIND for tera-scale mouse brain simulation with high fidelity. Using spatial gene expression as the input prompt, Tera-MIND achieves seamless generation of tera-scale mouse brain(s) under moderate hardware requirements, making it accessible for broader applications in both computational and biomedical research. In comparison to seven cutting-edge methods [22, 23, 42, 28, 12, 44, 46], which are developed upon either the GAN or diffusion framework, we thoroughly presented the overall improved performance achieved by Tera-MIND. Beyond its generation efficiency, Tera-MIND enables the comprehensive analysis of spatial gene-gene attention patterns, as exemplified for both GLUT and DOPA neuronal systems. These experimental analyses offer biologically relevant insights into the spatial organization and interaction of gene expression. Specifically, our findings are reported not only on the main mouse instance (Main), as primarily presented in [47], but also on two supporting mouse brains (Supp (m, f)) to ensure reproducibility and consistency. These results showcase the robustness of Tera-MIND in generating biologically accurate and spatially coherent brain morphologies guided by spatial molecular data.
We acknowledge the limitations of Tera-MIND. The modeling paradigm is fundamentally driven by the molecular-to-morphology spatial associations, which, while powerful, represent a simplified hypothesis of the complex neurological processes. Future refinements of this methodology could address this challenge by incorporating additional spatiotemporal data modalities, such as epigenomic, transcriptomic, proteomic, and (immuno-)histological information from diverse developmental and pathological contexts, along with adaptations for human tissue analysis. With these enhancements, Tera-MIND can further advance research into the spatial dynamics of gene expression and the cellular alterations associated with complex human (neurological) diseases [39].
In conclusion, the development of virtual replicas of entire biological organisms using GenAI, referred to as generative digital twins (GDTs [45]) and exemplified by Tera-MIND in this study, opens new avenues for high-throughput biomedical simulation with minimal ethical concerns. This methodology offers a cost-effective and scalable alternative for preclinical testing of therapeutic strategies. Lastly, the potentially wide applicability of GDTs on human samples aligns with the principles of animal welfare by supporting the replacement, reduction, and refinement (3R) of animal use in laboratory research, providing a promising algorithmic tool for ethically responsible biomedical innovation.
4 Methods
4.1 Spatial mRNA data as 3D ‘images’
Biological processes are inherently spatial [3]. The spatial molecular organization of the mammalian brain plays a decisive role in defining brain functionality and its dysfunction in disease. Recognizing the importance of spatial context, we leverage spatial mRNA readouts [47, 49] as 3D ‘images’, which are paired with corresponding 3D morphological bioimages. Unlike conventional data representations, such as 1D gene expression vector [8, 37], text description [6], graph-based structure [17, 1], or genomic sequence [27], our simple and vision-driven data format naturally maintains the 3D localization of molecular features, facilitating the holistic modeling of intrinsic spatial organizations that underpins brain function and dysfunction.
4.2 Spatial paired data processing
In prior studies [47, 49], three comprehensive and high-resolution brain atlases have been made publicly available  [^1] https://download.brainimagelibrary.org/aa/79/aa79b8ba5b3add56/. For each brain slice, there exists a large multiplex gene expression table and the associated morphological bioimages, as captured by DAPI-and PolyT-stained high-resolution WSIs ($0.108\mu m/\mathrm{pixel}$). Given the coordinate misalignment across available brain slices, we start the data processing by image registration. To maintain the original scale of raw data and for the sake of molecular data conversion, the registration-derived transformation matrix is required for our customized data processing. To this end, we employ the ABBA software [7] to perform the brain atlas registration, which allows us to access individual transformation matrices $\mathrm{S}$. By applying the inverse $\mathrm{S}^{-1}$ on the raw molecular and morphological data resp., we obtain the updated (large) gene expression table with transformed coordinates and the transformed WSI. For all brain slices, its transformed gene expression table is then converted to a sparse array, which has the same $73,216\times 105,984$ spatial resolution as the corresponding transformed WSI. Specifically, we employ the same processing pipeline on all the available atlases and obtain three sets of gene expression 3D images and associated WSIs. After excluding the brain slices with corrupted image quality and post-processing, we keep 50 slices of paired data for each brain atlas. Note that the processed results have been carefully examined by domain experts to ensure the registration quality. We refer interested readers to App. Fig. 8910 (a) for more details. After consecutively stacking 50 slices of gene expression arrays and WSIs resp., we obtain the whole tera-scale paired 3D ‘images’, from which we can collect a large amount of small data pairs (e.g., $128\times 128$) for running the model.
4.3 The proposed Tera-MIND
To faithfully capture the molecular-to-morphology spatial associations, we introduce a 3D gene-gene block that learns the spatial gene expression embeddings, which are subsequently integrated into the 3D gene-morph (gm) UNet to control the morphological reconstruction process (Fig. 1 (c, d)). Inspired by DIT [30, 50], we apply newly designed spatial attention blocks to our model architecture.
3D gene-gene (gg) self-attention. Instead of 1D conditional embeddings learned from textual descriptions, we use the 3D-gg self-attention layer to process the 3D gene expression array used as the conditional prompt. Let $\mathbf{g}\in\mathbb{R}^{n\times d}$ be a $n$-plex 3D gene expression array with the spatial dimension $d=\mathrm{width}\times\mathrm{height}\times z$, where $z$ denotes the number of stacked slices from which the gene expression is extracted. Then, we have
$\mathbf{Q_{g}}$ $=\mathsf{RMSNorm}(\mathbf{g}\mathbf{W_{q}}),\;\mathbf{K_{g}}=\mathsf{RMSNorm}(\mathbf{g}\mathbf{W_{k}}),$ (1)
$\mathbf{Attn_{g}}$ $=\underbrace{\mathsf{nn.Softmax}(\frac{\mathbf{Q_{g}}\mathbf{K_{g}^{T}}}{d})\mathbf{g}}_{\mathbf{Attn_{gg}}}\mathbf{W_{v}},\;\mathrm{where}\;\mathbf{V_{g}}=\mathbf{g}\mathbf{W_{v}}.$
Here, $\mathbf{W_{\{q,k,v\}}}\in\mathbb{R}^{d\times d}$ are learnable weights and $\mathbf{Attn_{gg}}$ is the derived spatial gene-gene attention reported in Fig. 3 and 5. After stepwise 3D convolutional upscaling on the $\mathbf{Attn_{g}}$, we have the sequential of 3D gene features $\mathbf{g}_{i}\in\mathbb{R}^{d_{i}\times n_{i}}$ for $i=1,\ldots,l$, with the increasing spatial resolutions by the magnitude of 2.
3D gene-morph (gm) cross-attention. To integrate the learned $\mathbf{g}_{i}$ into the reconstruction of morphological images by 3D-gm UNet, we inject selected $\mathbf{g}_{i}$ to the corresponding layer of encoder and decoder, which process morphological presentations $\mathbf{m}_{i}\in\mathbb{R}^{d_{i}\times c_{i}}$ at the same spatial resolution $d_{i}$. First, we obtain the spatial adaptive coefficients $\mathbf{Scale_{g_{i}}},\mathbf{Shift_{g_{i}}},\mathbf{Gate_{g_{i}}}\in\mathbb{R}^{d_{i}}$ and gene embedding $\mathbf{Embed_{g_{i}}}\in\mathbb{R}^{d_{i}}$ by inputting the $\mathbf{g}_{i}$ to the $\mathsf{AdaLN}$ [30, 50] block containing $\mathsf{nn.SiLU}$ and $\mathsf{nn.Linear}$ layers. Then, we have
$\mathbf{Ada_{m_{i}}}$ $=\mathsf{RMSNorm}(\mathbf{m_{i}})\cdot(\mathbf{Scale_{g_{i}}}+1)+\mathbf{Shift_{g_{i}}},$ (2)
$\mathbf{P_{m_{i}}}$ $=\mathbf{m_{i}}+\mathbf{Gate_{g_{i}}}\mathsf{nn.Softmax}(\frac{\mathbf{Q_{Ada_{m_{i}}}}\mathbf{K_{Ada_{m_{i}}}^{T}}}{c_{i}})\mathbf{V_{Embed_{g_{i}}}},$
where $\mathbf{Q_{Ada_{m_{i}}}}$ and $\mathbf{K_{Ada_{m_{i}}}}$ are query and key representations of $\mathbf{Ada_{m_{i}}}$, $\mathbf{V_{Embed_{g_{i}}}}$ is the value vector of $\mathbf{Embed_{g_{i}}}$. For both attention blocks, we use the Pytorch implementation of flash attention-2 [9] to improve computational efficiency and reduce memory cost.
Boundary-aware path. Following the denoising diffusion framework [15], we implement the forward process by gradually adding Gaussian noise to the 3D image patches $\mathbf{m}$ extracted from WSIs at timestep $t$, which gives us noisy image patches $\mathbf{m_{t}}$. Using 3D-gm UNet, we then parameterize the denoising function $\epsilon_{\theta}(\mathbf{m_{t}},t)$ to reverse the process. Similar as [12] and along with the standard denoising process, we supply an additional boundary-aware denoising path to impose the boundary consistency on the center-cropped patches. Therefore, our training objective is determined to be $\|\epsilon_{\theta}(\mathbf{m_{t}^{1}},t)-\epsilon_{1}\|^{2}+\|\epsilon_{\theta}(\mathbf{m_{t}^{2}},t)-\epsilon_{2}\|^{2}$, where $\mathbf{m_{t}^{1}}$ is the stack of small patches extracted from the input data and $\mathbf{m_{t}^{2}}$ is the center-cropped patch (See also Fig. 1 (c)). During the inference, we only use the boundary-aware path to generate image patches without stitching artifacts.
Model training. Among available mouse brain atlases, previous studies [47, 49] reported the primary results on the P56 female mouse brain (Main) including all the major brain regions (Fig. 1 (b)), while the additional P56 male and female mouse brains (Supp (m), (f)) that preserve most of the brain regions (Fig. 3 (a)) are used for supporting analyses. Following this experimental setup and to avoid overfitting, we report our primary generation result on the main mouse brain while taking the supp (m) and (f) mouse brains as the training data. Moreover, we filter out the genes that are present across three instances and obtain the 279-plex gene expression array for training (App. Fig. 8 (e)). In addition to the main results reported in Fig. 12, we conduct (unseen) generation experiments on the supporting mouse brains for the sake of reproducibility and consistency. Since supp (m) and (f) are processed using the same protocol with the identical $500$-plex gene list (App. Fig. 910 (e)), we take one ‘supp’ brain for training and the other unseen ‘supp’ brain for testing, and vice versa. Given the small distribution shift between supporting brain instances, this setup thus complements the challenging main experiment, which presents a large distribution gap between the training (supp (m, f)) and testing (main) data.
Model inference. After completing the model training, we run the tera-scale brain generation in a patch-wise manner. To ensure both computational feasibility and generation fidelity, we employ the Denoising Diffusion Implicit Models (DDIM) [36] approach for accelerated sampling and high-quality image synthesis. Given the significant GPU memory constraints associated with storing an entire mouse brain, the collection of generated patches is instead temporarily offloaded to the hard drive during each sampling step. These intermediate results are then utilized for the subsequent step. The patch-wise generation framework is inherently parallelizable, enabling efficient resource utilization. On a single NVIDIA A100 DGX system, the entire generation process for a tera-scale mouse brain can be completed within seven days. This demonstrates the scalability and practicality of Tera-MIND for handling large-scale 3D biomedical datasets.
4.4 Ablative studies
To assess the impact of key design choices in Tera-MIND, we discuss ablative results based on three critical factors: the number of slices used to extract 3D training data pairs, the spatial resolution of image patches, and the number of sampling steps. In line with the previous evaluations, both general metrics (PSNR and SSIM) and domain-specific metrics (Nuclear size and Cell number) are reported for a more comprehensive analysis.
For all three brain instances, optimal quantitative performance has been achieved when the slice number was set to 2. This is supported by the fact that neighboring slices are separated by a relatively large interval, ranging from $100\mu m$ to $200\mu m$ [47]. When compared to the neighboring pixels, which have a physical size of $0.108\mu m$ per pixel, the morphological context extracted from more neighboring slices is very likely saturated, as illustrated in Fig. 6 (a). Besides, in comparison to the 2D image training paradigm when the slice number equals 1, our proposed 3D modeling can further learn the cross-slice interaction patterns between neighboring slices. For instance, App. Fig. 7 highlights the consistent spatial Nr4a2-Th interaction patterns pinpointed in the very SN and VTA regions across three simulated mouse brains. These results substantiate the superiority of the proposed 3D modeling for the whole biological entity.
Additionally, we explore the effect of varying image patch resolution, ranging from $64\times 64$ to $256\times 256$ (Fig. 6 (b)). Here, we consider the entire patch size of the input data, rather than using an intermediate cropped patch size. Eventually, we determine that the resolution of $128\times 128$ yields the best performance. This choice aligns with the previous Patch-DM study[12], where the whole patch resolution was identified as 128 by default.
Finally, we examined the impact of the DDIM sampling step, weighing the trade-off between computational cost and performance. As shown in Fig. 6 (c), the optimal sampling step for Tera-MIND is determined to be 15. Beyond this point, additional steps lead to saturated improvements in performance, underscoring the cost-effectiveness of this choice.
Figure 6: The ablative results on the main and supp (m, f) mouse brains. a. Quantitative performance metrics evaluated across varying slice counts. b. Quantitative performance metrics assessed at different spatial resolutions. c. Quantitative performance metrics reported for varying sampling steps.
Data availability
All three mouse brain atlases, including the spatially resolved mRNA table and matched DAPI- and PolyT-stained WSIs for each brain slice, are accessible via https://download.brainimagelibrary.org/aa/79/aa79b8ba5b3add56/.
Code availability
The Python implementation used to train the model and reproduce the simulation results can be accessed via https://github.com/CTPLab/Tera-MIND.
Appendix
Figure 7: The gene-gene interaction visualization of the previous-current and current-next slices along posterior to anterior direction for main and supp mouse brains. a. The slice-wise visualization of Slc17a6 and Slc17a7 attention level (left) with regard to the DAPI- and PolyT-stained WSIs. b. The slice-wise visualization of Nr4a2 and Th attention level (left) with regard to the DAPI- and PolyT-stained WSIs.
[…]

### v3

@@ RESULTS
In dopamine-mediated neuronal signaling, Nr4a2 (Nurr1) is an important transcription factor that is expressed in several regions of the central nervous system including the olfactory bulb 31, the substantia nigra (SN) and ventral tegmental area (VTA). Nr4a2 plays a pivotal role in dopaminergic neurons by regulating several key genes (e.g., Th) involved in dopamine synthesis, storage, and release. Th encodes tyrosine hydroxylase, the rate-limiting enzyme in dopamine synthesis 31. Both Nr4a2 and Th genes are present in areas such as SN, VTA, and olfactory bulb of the mouse brain 32, 33.
The co-localization patterns have been correctly captured in spatial visualizations presented in Fig. 3 (e, left) and (f), which are subsequently used to derive spatial attention levels using self-attention mechanism (Fig. 3 (e, middle)). Reduced Nr4a2 and Th expression are associated with Parkinson’s disease 34, 35, a common neurodegenerative disorder with loss of dopaminergic neurons in the SN, leading to dopamine deficiency and motor function impairments 34, 35. This has also been observed in mice after deletion of Nr4a2 in mature dopaminergic neurons 36. Though the linear regression analysis of Fig. 3 (g) illustrates overall lower attention levels than the ones reported for GLUT system, we observe highly aggregated Nr4a2-Th attention signals in the olfactory bulb, SN and VTA (Bottom of midbrain label in Fig. 3 (h)), where both genes play a central role in regulating dopaminergic functions. As shown in Fig. 3 (e, right) and (h), the 3D gene-gene attention map uncovers a spatially aligned interaction between Nr4a2 and Th that reinforces their roles in the molecular regulation of dopaminergic pathways and cellular anatomical structures.
Tera-MIND achieved reproducible and robust results on three tera-scale mouse brains.
In addition to the main mouse brain discussed above, two additional P56 adult mouse brains – one male and one female – were profiled in previous studies 1, 2 for supporting analyses, as illustrated in Fig. 4 (a) (middle and right). To rigorously evaluate the reproducibility and robustness of our findings, we extend the same simulation experiments conducted on the main mouse brain to both supporting mouse brains (supp (m) and supp (f)).
Morphology generation: As illustrated in Fig. 4 (a, b), the side-by-side visual comparisons between our generated and GT results demonstrate convincing reconstructions of brain morphology on all three tera-scale mouse brains. Consistent with the main generation results (Fig. 1 (b) and the left plots of Fig. 4 (a, b)), our supporting experiments highlight competitive and reliable 3D and slice-wise image generation quality, as elaborated by the holistic and multi-scale visualization shown in the middle and right plots of Fig. 4 (a, b). For quantitative evaluation, Tera-MIND is benchmarked against seven SOTA methods COCO-GAN 20, InfinityGAN 21, MS-PIE 22, SST-editing 37, IST-editing 25, SinFusion 23, and Patch-DM 24. Complementary to PSNR and SSIM scores, we incorporate the spatial Fréchet Inception Distance (sFID) 38, which prioritizes spatially relevant evaluations by rewarding image distributions that preserve coherent spatial structures. Across all three brain instances, Tera-MIND consistently outperforms competing methods, in terms of achieving better PSNR, SSIM, and sFID scores (Fig. 4 (c)). When evaluated on domain-specific metrics such as cell number and nuclear size (Fig. 4 (d, e)), Tera-MIND yields minimal distributional discrepancies relative to GT. This highlights its capacity to maintain high-fidelity not only in structural reconstruction but in the reproduction of biologically relevant features.
Pathway identification: In the context of the GLUT pathway, Fig. 5 (a, b) present a highly consistent and spatially comparable localization of Slc17a6 and Slc17a7 expression patterns across two independent supporting instances. These expression maps are accurately registered in 3D space and align well with the ones from our main results and previous studies 1, 2. Similarly as observed in the main results, Fig. 5 (c) and (d) illustrate that Tera-MIND is capable of learning strong and heterogeneous Slc17a6-Slc17a7 attention levels driven by the positively correlated gene expression levels. This suggests their synergistic and collective roles in regulating neurotransmission and cellular morphology across overlapped regions. Regarding the DOPA pathway, despite the absence of the olfactory bulb in the raw data of the supporting mouse brains, the gene expression patterns of Nr4a2 and Th exhibit a comparable emergence in the SN and VTA regions (See also Fig. 5 (e, f)). Consistent with main results, the linear regression analysis of Fig. 5 demonstrates the comparable correlation between gene expression and attention levels. Moreover, we witness reproducible and clear identification of Nr4a2-Th attention levels derived from Tera-MIND in the same regions for both supporting mouse brains (Fig. 5 (e, h)). These results confirm a consistent expression pattern of critical genes associated with dopaminergic function in these specific regions. Please see also the side-by-side video visualizations for more details.
Tera-MIND facilitated the translational application on unseen human brain samples.
Following the translational aims of mouse brain studies in neuroscience, we conduct generalization experiments in which Tera-MIND was trained exclusively on mouse brain data (Main setting) and tested on the human brain healthy (Fig. 6 (a)) and glioblastoma (Fig. 6 (b)) samples  [^1] https://www.10xgenomics.com/datasets/xenium-human-brain-preview-data-1-standard. As shown in the gene panel (c) of Fig. 6, we firstly determine the targeted genes that are present in both mouse brain atlases and human brain samples, which are then fed into the Tera-MIND model as the conditional input. After training with the patch-wise paired data from mouse brain atlases, we run the WSI generation using the gene expression arrays profiled on the healthy and glioblastoma human samples. Both qualitative (Fig. 6 (a, b)) and quantitative evaluations (Fig. 6 (d)) confirm that the performance of morphological simulation is generalizable from mouse to human samples, in terms of well-preserved tissue architecture and meaningful cellular correspondence to GT. Notably, Tera-MIND reported better PSNR and SSIM for the healthy brain sample than in the setting of Glioblastoma, suggesting a mild cross-species translational gap under non-physiological conditions.
When examining the gene expression patterns of GLUT pathway between healthy and glioblastoma cases, Fig. 6 (a) and (b) reveal distinct spatial distributions: Slc17a7 exhibits higher and more heterogeneous expression levels across the glioblastoma sample compared to the healthy sample, resulting in pronounced interaction signals between Slc17a6 and Slc17a7. Conversely, both gene expression and their attention levels for the healthy sample remain homogeneous. This phenomenon can be attributed to the observation that glioma cells often release glutamate, leading to excitotoxic neuronal death and creating space for tumor expansion 39.
DISCUSSION
In this study, we proposed a novel patch-based diffusion model Tera-MIND for high-fidelity simulation of mammalian brains. For three tera-scale mouse brain atlases, Tera-MIND achieved reproducible generation results and consistently identified spatial molecular interactions within GLUT and DOPA pathways. Beyond the comprehensive evaluation on murine brains, we established the in-silico translational applicability of Tera-MIND by generalizing to previously unseen human brain samples, including both healthy and glioblastoma tissues. Methodologically, Tera-MIND offers an efficient and scalable generative framework for modeling whole (animal) organisms. From the viewpoint of clinical application, Tera-MIND thereby provides a novel in-silico approach for comparative pathology 40, which enables a systematic and side-by-side analysis of murine and human tissues to assess the validity and translational relevance of animal models. Furthermore, Tera-MIND inherently facilitates in-silico interventions on targeted genes of interest that drive (disease) morphological transitions 25. This capability allows for the exploration of ‘what if’ causal questions for enhancing clinical treatment and diagnostics 8. Given the scope of this study, we plan to address the topic of simulated intervention in future work.
Limitations of the study
In addition, we acknowledge the limitations of Tera-MIND. The modeling paradigm is fundamentally driven by the molecular-to-morphology spatial associations, which, while powerful, represent a simplified hypothesis of the complex neurological processes. With the rapid emergence of diverse spatial omics technologies 41, future work will incorporate orthogonal modalities, such as spatial proteomics and epigenomics, to provide additional layers of validation and further distinguish true functional spatial relationships.
In conclusion, the development of virtual replicas of entire biological organisms using GenAI, referred to as generative digital twins (GDTs 8) and exemplified by Tera-MIND in this study, opens new avenues for high-throughput biomedical simulation with minimal ethical concerns. This methodology offers a cost-effective and scalable alternative for preclinical testing of therapeutic strategies. Lastly, the potentially wide applicability of GDTs on human samples aligns with the principles of animal welfare by supporting the replacement, reduction, and refinement (3R) of animal use in laboratory research, providing a promising algorithmic tool for ethically responsible biomedical innovation.
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Jiqing Wu (Jiqing.Wu@unibas.ch).
Data and code availability
  • The processed mouse brain data has been deposited at brain image library under the database identifier ace-lot-now and are publicly available as of the date of publication. All other data reported in this paper will be shared by the lead contact upon request.
  • All original code has been deposited at Zenodo under the DOI 10.5281/zenodo.14826874 and is publicly available as of the date of publication.
  • Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
ACKNOWLEDGMENTS
This research project is mainly supported through the core professorship funding of Prof. Dr. med. Viktor Koelzer, Medical Co-Director and Professor of Experimental Pathology, University Hospital of Basel.
AUTHOR CONTRIBUTIONS
Conceptualization, J.W. and V.H.K.; methodology, J.W.; investigation, J.W., I.B., V.H.K.; writing-–original draft, J.W., V.H.K.; writing-–review & editing, J.W., I.B., V.H.K., Y.L., E.K.; funding acquisition, V.H.K.; supervision, V.H.K..
DECLARATION OF INTERESTS
V.H.K. reports being an invited speaker for Sharing Progress in Cancer Care (SPCC) and Indica Labs; advisory board of Takeda; and sponsored research agreements with Roche and IAG, all unrelated to the current study. V.H.K. is a participant of patent applications in computational pathology outside ofthe submitted work.
DECLARATION OF GENERATIVE AI AND AI-ASSISTED TECHNOLOGIES
During the preparation of this work, the author(s) used ChatGPT in order to polish the content. After using this tool or service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
SUPPLEMENTAL INFORMATION INDEX
Figures S1-S5 and their legends in a PDF
MAIN FIGURE TITLES AND LEGENDS
Figure 1: Overview of the Tera-MIND architecture and workflows for virtual organ simulation (Main result). a. The 3D reference map of major brain regions (left) alongside their corresponding labels (right) for guiding the navigation of brain structures. b. The 3D visualization of the ground-truth (GT, left) and generated (right) mouse brain, as captured by a sequential stack of DAPI and PolyT-stained WSIs at the scale of $0.77\times 10^{12}$ voxels. The bottom part of generated and GT results overlay brain region maps, aiding in the navigation of the complex spatial brain architecture. c. The conceptual illustration of patch-based diffusion model training. Here, we train Tera-MIND with noisy DAPI and PolyT bioimage patches so that the model learns to output clean patches guided by paired spatial mRNA array. Apart from the standard reconstruction path (black arrows), we supply the boundary-aware path (gray arrows) to impose boundary consistency between generated neighboring patches. d. The conceptual illustration of patch-wise generating virtual mouse brain using Tera-MIND. At this stage, we only run through the boundary-aware path to seamlessly generate the tera-scale mouse brain.
Figure 2: The visual and quantitative comparison of mouse brain generation (Main result). a. The 3D map of major brain regions (left), GT (middle), and the generated mouse brain (right) of the atlas used in the primary analysis 1. b. The cell- and region-level visualization of GT mouse brain WSIs in selected brain slices. For (a) and (b), the 3D/2D reference maps of brain regions are overlaid to aid in the navigation of brain structures. c. The cell- and region-level visualization of generated mouse brain WSIs. d. The visual comparison of exemplar cellular regions ($1024\times 1024$) between Tera-MIND and state-of-the-art (SOTA) patch-based models. From left to right, the image stack displays individual DAPI, PolyT, merged channels, and nuclear masks obtained using Cellpose 42. e. Quantitative comparison of image generation quality between Tera-MIND and SOTA methods. This includes mean and standard deviation scores (error bar) for widely-used image quality metrics, such as PSNR and SSIM, alongside domain-specific scores like patch-based nuclear size and cell count. For a more in-depth analysis, we encourage readers to zoom in on the visual results for clearer comparisons.
@@ ACKNOWLEDGMENTS
Figure 3: The visualization of gene-gene interactions for GLUT and DOPA neuronal systems (Main result). a. The 3D visualization of Slc17a6 and Slc17a7 gene expression (left), the conceptual illustration of 3D gene-gene attention layer (middle), and the 3D visualization of Slc17a6-Slc17a7 attention (right). b. The WSI visualization of Slc17a6 and Slc17a7 gene expression (expr.). c. The linear regression analysis of gene expr. and attention level for both Slc17a6 and Slc17a7. d. The WSI visualization of Slc17a6-Slc17a7 attention level. e. The 3D visualization of Nr4a2 and Th gene expression (left), the conceptual illustration of 3D gene-gene attention layer (middle), and the 3D visualization of Nr4a2-Th attention (right). f. The WSI visualization of Nr4a2 and Th gene expression (expr.). g. The trending line showing the relationship between gene expr. versus attention level for both Nr4a2 and Th. h. The WSI visualization of Nr4a2-Th attention level. Here, both SN and VTA are located towards the bottom of midbrain label. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for both GLUT and DOPA pathways.
@@
Figure 5: The visualization of gene-gene interaction for GLUT and DOPA neuronal systems (Supp (m) and (f) results). a. The 3D visualization of Slc17a6 and Slc17a7 gene expression (left) and the 3D visualization of Slc17a6-Slc17a7 attention (right). b. The WSI visualization of Slc17a6 and Slc17a7 gene expression (expr.). c. The linear regression analysis of gene expr. and attention level for both Slc17a6 and Slc17a7. d. The WSI visualization of Slc17a6-Slc17a7 attention level. e. The 3D visualization of Nr4a2 and Th gene expression (left) and the 3D visualization of Nr4a2-Th attention (right). f. The WSI visualization of Nr4a2 and Th gene expression (expr.). g. The linear regression analysis of gene expr. and attention level for both Nr4a2 and Th. h. The WSI visualization of Nr4a2-Th attention level. Here, both SN and VTA are located towards the bottom of midbrain label. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for both GLUT and DOPA pathways.
Figure 6: The comprehensive generation, quantification, and gene expression statistics for the human brain experiments. a. The side-by-side comparison between the generated (Gen) and ground-truth (GT) WSIs of the human brain (Healthy), with a $27136\times 36864$ spatial resolution. The visualization of Slc17a6 and Slc17a7 gene expression and their spatial attention. b. The side-by-side comparison between the Gen and GT WSIs of the human brain (Glioblastoma), with a $23808\times 39680$ spatial resolution. The visualization of Slc17a6 and Slc17a7 gene expression and their spatial attention. c. The side-by-side comparison of gene expression level (mean) for the human (Healthy) and human (Glioblastoma). d. The quantitative comparison of generation quality for the human (Healthy) and human (Glioblastoma), as measured by PSNR, SSIM, nuclear size, and cell number. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for the GLUT pathway.